# Improving the (approximate) sequential probability ratio test by avoiding overshoot

### v1

@@ 6.1 Testing a forecaster
The classical SPRT by Wald [34, 35] assumes that the null hypothesis and the alternative hypothesis are specified in advance in the form of simple distributions $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$, respectively. However, in Section 3.2 we have seen that the SPRT remains even valid if we replace the fixed alternative $\mathbb{P}_{1}$ with a predictable plugin $\mathbb{P}_{1}^{t}$ that depends on the data $X_{1},\ldots,X_{t-1}\in\mathcal{X}$ observed so far. In general, the null distribution $\mathbb{P}_{0}^{t}$ and the alternative distribution $\mathbb{P}_{1}^{t}$ considered at each step $t\in\mathbb{N}$ can be chosen freely based on the previous data $X_{1},\ldots,X_{t-1}$ and some additional information $Z_{1},\ldots,Z_{t}\in\mathcal{Z}$. Shafer [22] formulated this as a forecaster-sceptic testing protocol, where the forecaster announces a probability distribution $\mathbb{P}_{0}^{t}$ at each step $t$ and a sceptic tests the forecaster by specifying an alternative $\mathbb{P}_{1}^{t}$. For example, one can think of the forecaster as a weather forecaster, which is tried to be invalidated by a sceptic. We summarized this general testing protocol in Algorithm 2.
The extension of our boosting technique (Algorithm 1) to the general testing problem in Algorithm 2 is straightforward, since at each step $t$ we only require the current martingale value $M_{t-1}=\Lambda_{t-1}$ and the distribution of the next factor $L_{t}=\frac{\mathbb{P}_{1}^{t}(X_{t})}{\mathbb{P}_{0}^{t}(X_{t})}$ under the null hypothesis $\mathbb{P}_{0}^{t}$.
Algorithm 2 SPRT in a general forecaster-sceptic testing protocol
@@
5:     Sceptic announces a probability distribution $\mathbb{P}_{1}^{t}$ on $\mathcal{X}$
@@
7:     $\Lambda_{t}=\Lambda_{t-1}\cdot\frac{\mathbb{P}_{1}^{t}(X_{t})}{\mathbb{P}_{0}^{t}(X_{t})}$
@@ 6.2 Other approaches to avoid Wald’s approximated thresholds
The problem of “overshoot” in the SPRT was already noted by Wald [34, 35] and examined in many previous works. One line of work approximates the exact type I and type II error of SPRTs based on Brownian motion and renewal theory [23, 25, 3, 12, 43, 15]. While the resulting approximations are asymptotically correct, they are very difficult to compute and therefore only used rarely in practice [23, 25]. For example, we have already seen that the available R packages for performing an SPRT [27, 2] rely on Wald’s approximated thresholds and that most applied studies also use them (e.g. [26, 13, 4, 9]). Another option is to truncate the SPRT by prespecifying a maximum number of observations. For simple distributions of the likelihood ratio and sufficiently small maximum numbers of observations it is then possible to calculate the type I and type II error for given thresholds exactly [1], which can be used to find appropriate bounds for the SPRT. Such an approach is frequently employed in group sequential trials [14, 39]. For more complex distributions one could approximate the error probabilities using Monte-Carlo sampling. This often leads to considerably computational effort, which sometimes can be reduced using importance sampling [24]. A limitation these three approaches have in common is that the used test statistics including their distributions must be specified in advance. This is not necessarily the case in the general forecaster-sceptic testing protocol described in the previous section. Hence, while our boosting approach is applicable to such complex and adaptive testing protocols, all of the aforementioned methods fail.
@@ 7 Summary
In this paper, we introduced a general approach to uniformly improve nonnegative (super)martingale based sequential tests by avoiding an overshoot at the upper threshold $1/\alpha$ (where $\alpha$ is the desired type-I error). We illustrated its application by the power-one SPRT with one-parameter exponential families for one-sided null and alternative hypotheses. We also showed how it can be applied when a lower threshold is specified, that determines an early stopping time for accepting the null hypothesis. With this we obtain an alternative to the two-sided SPRT with Wald’s approximated thresholds, where our alternative provides valid type I error control and leads to similar type II error as Wald’s proposal while often reducing the sample size. Our sequential tests can also be used to tighten confidence sequences and be extended to nontrivial settings like sampling without replacement and conformal martingales.
@@ Acknowledgments
LF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 281474342/GRK2224/2. AR was funded by NSF grant DMS-2310718.

### v2

@@ 6.1 Testing a forecaster
The classical SPRT by Wald [36, 37] assumes that the null hypothesis and the alternative hypothesis are specified in advance in the form of simple distributions $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$, respectively. However, in Section 3.2 we have seen that the SPRT remains even valid if we replace the fixed alternative $\mathbb{P}_{1}$ with a predictable plugin $\mathbb{P}_{1}^{t}$ that depends on the data $X_{1},\ldots,X_{t-1}\in\mathcal{X}$ observed so far. In general, the null distribution $\mathbb{P}_{0}^{t}$ and the alternative distribution $\mathbb{P}_{1}^{t}$ considered at each step $t\in\mathbb{N}$ can be chosen freely based on the previous data $X_{1},\ldots,X_{t-1}$ and some additional information $Z_{1},\ldots,Z_{t}\in\mathcal{Z}$. Shafer [24] formulated this as a forecaster-skeptic testing protocol, where the forecaster announces a probability distribution $\mathbb{P}_{0}^{t}$ at each step $t$ and a skeptic tests the forecaster by specifying an alternative $\mathbb{P}_{1}^{t}$. For example, one can think of the forecaster as a weather forecaster, which is tried to be invalidated by a skeptic. We summarized this general testing protocol in Algorithm 2.
The extension of our boosting technique (Algorithm 1) to the general testing problem in Algorithm 2 is straightforward, since at each step $t$ we only require the current martingale value $M_{t-1}=\Lambda_{t-1}$ and the distribution of the next factor $L_{t}=\frac{p_{1}^{t}(X_{t})}{p_{0}^{t}(X_{t})}$ under the null hypothesis $\mathbb{P}_{0}^{t}$, where $p_{0}^{t}$ and $p_{1}^{t}$ are the densities of $\mathbb{P}_{0}^{t}$ and $\mathbb{P}_{1}^{t}$, respectively.
Algorithm 2 SPRT in a general forecaster-skeptic testing protocol
@@
5:     Skeptic announces a probability distribution $\mathbb{P}_{1}^{t}$ on $\mathcal{X}$
@@
7:     $\Lambda_{t}=\Lambda_{t-1}\cdot\frac{p_{1}^{t}(X_{t})}{p_{0}^{t}(X_{t})}$
@@ 6.2 Other approaches to avoid Wald’s approximated thresholds
The problem of “overshoot” in the SPRT was already noted by Wald [36, 37] and examined in many previous works. One line of work approximates the exact type I and type II error of SPRTs based on Brownian motion and renewal theory [25, 27, 3, 13, 45, 17]. While the resulting approximations are asymptotically correct, they are very difficult to compute and therefore only used rarely in practice [25, 27]. For example, we have already seen that the available R packages for performing an SPRT [29, 2] rely on Wald’s approximated thresholds and that most applied studies also use them (e.g. [28, 14, 4, 10]). Another option is to truncate the SPRT by prespecifying a maximum number of observations. For simple distributions of the likelihood ratio and sufficiently small maximum numbers of observations it is then possible to calculate the type I and type II error for given thresholds exactly [1], which can be used to find appropriate bounds for the SPRT. Such an approach is frequently employed in group sequential trials [15, 41]. For more complex distributions one could approximate the error probabilities using Monte-Carlo sampling. This often leads to considerably computational effort, which sometimes can be reduced using importance sampling [26]. A limitation these three approaches have in common is that the used test statistics including their distributions must be specified in advance. This is not necessarily the case in the general forecaster-skeptic testing protocol described in the previous section. Hence, while our boosting approach is applicable to such complex and adaptive testing protocols, all of the aforementioned methods fail.
@@ 7 Summary
In this paper, we introduced a general approach to uniformly improve nonnegative (super)martingale based sequential tests by avoiding an overshoot at the upper threshold $1/\alpha$ (where $\alpha$ is the desired type-I error). We illustrated its application by the power-one SPRT with one-parameter exponential families for one-sided null and alternative hypotheses. We also showed how it can be applied when a lower threshold is specified, that determines an early stopping time for accepting the null hypothesis. With this we obtain an alternative to the two-sided SPRT with Wald’s approximated thresholds, where our alternative provides valid type I and type II error control while often reducing the sample size. Our sequential tests can also be used to tighten confidence sequences and be extended to nontrivial settings like sampling without replacement and conformal martingales.
@@ Acknowledgments
LF acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project number 281474342/GRK2224/2. AR was funded by NSF grant DMS-2310718. We are grateful to Wouter M. Koolen for pointing out the possibility of using an inverted SPRT to control the type II error, and to Nick Koning for mentioning the log-optimality of boosting factors. In addition, we would like to thank Yajun Mei and Qunzhi Xu for a useful discussion.
Appendix A Additional simulation results
In this section, we provide some additional simulation result. In Figure 6 we repeated the simulated experiment for simple hypotheses from Figure 1 but with $\alpha=0.01$ instead of $\alpha=0.05$. Obviously, the required sample size is larger for $\alpha=0.01$ than for $\alpha=0.05$. It can also be seen that the gain in percentage of sample size is smaller than for $\alpha=0.05$, however, due to the larger total sample size the absolute gain in sample size remains similar.
Figure 6: Comparison of sample size and type I error between the SPRT and the boosted SPRT in a simple Gaussian testing setup for different signal strengths and $\alpha=0.05$. The boosted SPRT saves $2-8\%$ of sample size compared to the SPRT and exhausts the significance level. The SPRT is becoming increasingly conservative for stronger signals.
In Figure 7, we repeated the two-sided experiment from Figure 5 for $\alpha=0.01$. Again, the boosted SPRT saves a substantial number of samples compared to Wald’s two-sided SPRT with approximated thresholds and nearly exhausts the type I and type II error. Interestingly, the percentage of saved samples approximately increases with $\beta$ for $\alpha=0.01$, while it decreases for $\alpha=0.05$.
Figure 7: Comparison of sample size, type II and type I error probability between the SPRT with Wald’s approximated thresholds and the boosted SPRT in a simple Gaussian testing setup for different desired type II error probabilities ($\beta$) and $\alpha=0.01$. The boosted method provably controls the type I and type II error, while the SPRT with Wald’s approximated thresholds does not. Furthermore, the boosted SPRT saves sample size and nearly exhausts the error probabilities.